OpenStudy (anonymous):

Use substitution to determine if the given ordered pair is an element of the solution set for the system of equations. (-2, 2) 5y - 5x = 10 3x + 10 = 2y Solution Not a solution

OpenStudy (lucaz):

as the problem says, substitute x and y and see if the equation works

OpenStudy (anonymous):

i did i got for the 1st one, 10+10=10 and the 2nd .. -6+10=4 soo, is that not a solution?

OpenStudy (anonymous):

do both answers have to be the same to be solution?

OpenStudy (lucaz):

1st is 5(2)-5(-2)=10

OpenStudy (turingtest):

both equations do not have to be the same, they just both have to be true

OpenStudy (acxbox22):

not a solution because the ordered pair does not work for both equations

OpenStudy (anonymous):

See thats what i thought

OpenStudy (turingtest):

@acxbox22 would you care to at least demonstrate that assessment?

OpenStudy (anonymous):

So its a solution?

OpenStudy (lucaz):

it's a solution if the x and y values make both equations true, you substitute the values and see if that's the case

OpenStudy (anonymous):

Yal all have me confused lol, i thought it was not a solution

OpenStudy (lucaz):

5(2)-5(-2)=10 calculate this

OpenStudy (anonymous):

its 10+10=10

OpenStudy (lucaz):

10+10 is equal to 20 not 10 so this doesnt work

OpenStudy (lucaz):

(-2,2) would be a solution only if it worked for both equations, you already know this is not true